# Quantum Wiretap Channel Coding Assisted by Noisy Correlation

### v1

@@ I Introduction
Secure communication over a classical channel with an eavesdropper was first introduced by Wyner [32]. The secrecy capacity for quantum wiretap channels has been determined in [16] and [10]. Chen et al. [12] demonstrated that correlation between Alice and Bob is able to help secure message transmission over a classical wiretap channel. Inspired by these results, we analyze secure communication when the channel users (Alice, Bob and Eve) share a general correlated states, which are “given by nature”, and assumed to be known to all parties. In this work, we show a way to increase the secure capacity of quantum channels by the assistance of such noisy correlation. Our proof works by considering the shared resource as a component of the channel. Technically, we employ quantum wiretap channel codes, and to incorporate the correlated resource a variant of Gel’fand-Pinsker coding as in [13] and [2], which has been used similarly in secure “writing on dirty paper” codes [14].
@@ II Communication scenario
Before starting, we review some basic notions from quantum information theory. Let $\rho_{1}$ and $\rho_{2}$ be Hermitian operators on a finite-dimensional complex Hilbert space $A$. We say $\rho_{1}\geq\rho_{2}$ and $\rho_{2}\leq\rho_{1}$ if $\rho_{1}-\rho_{2}$ is positive semidefinite. We denote the set of density operators (states) on $A$ by $\mathcal{S}(A):=\{\rho\in\mathcal{L}(A):\rho\geq 0,\ \operatorname{Tr}(\rho)=1\}$, where $\mathcal{L}({A})$ is the set of linear operators on ${A}$. For finite-dimensional complex Hilbert spaces $A$ and $B$, a quantum channel is a linear, completely positive and trace preserving (cptp) map $\mathcal{N}:\mathcal{L}(A)\rightarrow\mathcal{L}(B)$, that acts a $\mathcal{L}(A)\ni\rho\mapsto\mathcal{N}\rho)\in\mathcal{L}(B)$, which accepts input quantum states in $\mathcal{S}(A)$ and produces output quantum states in $\mathcal{S}(B)$. We shall often simplify the notation of a quantum channel as $\mathcal{N}:A\rightarrow B$. Likewise, we will often suppress the tensor product sign in $BE=B\otimes E$ and $A^{n}=A^{\otimes n}$, if there is no danger of confusion.
@@
$\frac{1}{M}\left\|\operatorname{Tr}_{B^{n}{B^{\prime}}^{n}}(\mathcal{N}^{\otimes n}\!\circ\mathcal{E}_{m})\zeta^{\otimes n}-\sigma^{E^{n}{E^{\prime}}^{n}}\right\|_{1}$ $\leq\mu.$ (2)
See Fig. 1 for the communication diagramme.
@@ III Preliminaries
We shall nned two commonly used measures of distance and similarity of quantum states. The fidelity of two quantum states $\rho$ and $\sigma$ is defined as $F(\rho,\sigma)=\|\sqrt{\rho}\sqrt{\sigma}\|_{1}=\operatorname{Tr}\sqrt{\sqrt{\rho}\sigma\sqrt{\rho}}$, where $\|X\|_{1}=\operatorname{Tr}\sqrt{XX^{\dagger}}$ is the trace norm. Then, according to Fuchs and van de Graff,
@@ IV Main result
Looking again at the definitions and the communication diagramme in Fig. 1, we note that not just a code but any given modulations $\mathcal{E}_{u}$ ($u\in\mathcal{U}$) give rise to a cq-wiretap channel $\mathcal{M}:\mathcal{U}\rightarrow(BB^{\prime})^{n}(EE^{\prime})^{n}$, mapping $u\in\mathcal{U}$ to $\mathcal{M}(u)=(\mathcal{N}^{\otimes n}\!\circ\mathcal{E}_{u})\zeta^{\otimes n}$. We can thus apply the coding theorem and weak converse from [10] to get the following expression for the assisted private capacity:
@@
V Conclusion
We have presented a coding strategy for the quantum wiretap channel assisted by a general quantum correlation, which shows a systematic improvement of the private communication rate over the unassisted private capacity of the channel due to the exploitation of the correlation. Our Theorem 5 alos directly generalises the main result of [12], since classical channels are a special case of cptp maps and a classical correlation between random variables $X$, $Y$ and $Z$ is naturally represented by the diagonal state $\zeta^{A^{\prime}B^{\prime}E^{\prime}}=\sum_{xyz}P_{XYZ}(xyz)\left|x\right\rangle\!\left\langle x\right|^{A^{\prime}}\otimes\left|y\right\rangle\!\left\langle y\right|^{B^{\prime}}\otimes\left|z\right\rangle\!\left\langle z\right|^{E^{\prime}}$. As the proposed code uses a method derived from the (quantum) Gel’fand-Pinsker wiretap channel with side information, it is also always at least as good as, or superior to, the “trivial” incorporation of the correlation resource into an augmented wiretap channel. However, due to the general need for regularisation, both our main result (Theorem 5) and the “trivial” achievable rate [Eq. (5)] lead to multi-letter formulas for the same quantity $P(\mathcal{N},\zeta)$.
@@ V Conclusion
Now this achievability bound and expression for the assisted capacity correspond exactly to Eq. (3), and interestingly it provides and instance where Theorem 5 is better, at least on the single-letter level (naturally, in the regularisation both expression result in the same number). Namely, in [21] a certain duality relation was observed between the dense coding advantage and the so-called entanglement of purification $E_{P}$ [27], which extends to the regularisations of the two quantities:
@@
It remains unknown whether $\Delta=\Delta^{(\infty)}$ and $E_{P}=E_{P}^{(\infty)}$, but in [11] numerical evidence to the contrary was provided for the entanglement of purification Showing $E_{P}(\rho)>E_{P}^{(\infty)}(\rho)$ for a certain state in [11] required calculating the l.h.s. numerically (which is the evidence for which mathematical proof is lacking), and finding a new upper bound for the r.h.s. via an asymptotic protocol using the covering lemma. Concretely, it was shown for an ensemble $\{q(u),\rho_{u}^{CD}\}$ with $\rho=\sum_{u}q(u)\rho_{u}$ that
@@
where $\gamma^{UCD}=\sum_{u}q(u)\left|u\right\rangle\!\left\langle u\right|^{U}\otimes\rho_{u}^{CD}$ is the cq-state of the ensemble. If we us the duality relation from [21] to translate this protocol to $\Delta^{(\infty)}$, we can see that it corresponds to the achievability bound from Theorem 5 in the case $\mathcal{N}=\operatorname{id}_{A}$ and $\zeta^{A^{\prime}B^{\prime}}$.
@@ Acknowledgments
MLC was supported by the German Research Foundation (DFG) under the Walter Benjamin Fellowship CA-2779/1-1. AW is supported by the European Commission QuantERA grant ExTRaQT (Spanish MICIN project PCI2022-132965), by the Spanish MICIN (project PID2022-141283NB-I00) with the support of FEDER funds, the Spanish MICIN with funding from European Union NextGenerationEU (PRTR-C17.I1) and the Generalitat de Catalunya, by the Alexander von Humboldt Foundation, and the Institute for Advanced Study of the Technical University Munich.

### v2

@@ I Introduction
Secure communication over a classical channel with an eavesdropper was first introduced by Wyner [34]. The secrecy capacity for quantum wiretap channels has been determined in [16] and [10]. Chen et al. [12] demonstrated that correlation between Alice and Bob is able to help secure message transmission over a classical wiretap channel. Inspired by these results, we analyze secure communication when the channel users (Alice, Bob and Eve) share a general correlated state, which is “given by nature”, and assumed to be known to all parties. In this work, we show a way to increase the secure capacity of quantum channels by the assistance of such noisy correlation. Our proof works by considering the shared resource as a component of the channel. Technically, we employ quantum wiretap channel codes, and incorporate the correlated resource via a variant of Gel’fand-Pinsker coding as in [13] and [2]. This has been used similarly in secure “writing on dirty paper” codes [14].
@@
We would like to mention another coding protocol, which has appeared previously in the literature, showing a similar result [31]: this paper determined the entanglement generation capacity of a quantum channel with shared quantum states as a resource. The proof of the achievability part uses a similar approach to the dirty paper codes, although our construction cannot be reduced to theirs, in the same way as secret key generation is more general than entanglement generation [21]. See also [16] for the first description of codes achieving the entanglement generation capacity based on wiretap codes.
@@ II Communication scenario
Before starting, we review some basic notions from quantum information theory. Let $\rho_{1}$ and $\rho_{2}$ be Hermitian operators on a finite-dimensional complex Hilbert space $A$. We say $\rho_{1}\geq\rho_{2}$ and $\rho_{2}\leq\rho_{1}$ if $\rho_{1}-\rho_{2}$ is positive semidefinite. We denote the set of density operators (states) on $A$ by $\mathcal{S}(A):=\{\rho\in\mathcal{L}(A):\rho\geq 0,\ \operatorname{Tr}(\rho)=1\}$, where $\mathcal{L}({A})$ is the set of linear operators on ${A}$. For finite-dimensional complex Hilbert spaces $A$ and $B$, a quantum channel is a linear, completely positive and trace preserving (cptp) map $\mathcal{N}:\mathcal{L}(A)\rightarrow\mathcal{L}(B)$, that acts a $\mathcal{L}(A)\ni\rho\mapsto\mathcal{N}(\rho)\in\mathcal{L}(B)$, which accepts input quantum states in $\mathcal{S}(A)$ and produces output quantum states in $\mathcal{S}(B)$. We shall often simplify the notation of a quantum channel as $\mathcal{N}:A\rightarrow B$. Likewise, we will often suppress the tensor product sign in $BE=B\otimes E$ and $A^{n}=A^{\otimes n}$, if there is no danger of confusion.
@@
$\frac{1}{M}$ $\sum_{m}\left\|\operatorname{Tr}_{B^{n}{B^{\prime}}^{n}}(\mathcal{N}^{\otimes n}\!\circ\mathcal{E}_{m})\zeta^{\otimes n}-\sigma^{E^{n}{E^{\prime}}^{n}}\right\|_{1}\leq\mu.$ (2)
See Fig. 1 for the communication diagram.
@@ III Preliminaries
We shall need two commonly used measures of distance and similarity of quantum states. The fidelity of two quantum states $\rho$ and $\sigma$ is defined as $F(\rho,\sigma)=\|\sqrt{\rho}\sqrt{\sigma}\|_{1}=\operatorname{Tr}\sqrt{\sqrt{\rho}\sigma\sqrt{\rho}}$, where $\|X\|_{1}=\operatorname{Tr}\sqrt{XX^{\dagger}}$ is the trace norm. Then, according to Fuchs and van de Graff,
@@ IV Main result
Looking again at the definitions and the communication diagram in Fig. 1, we note that not just a code but any given modulations $\mathcal{E}_{u}$ ($u\in\mathcal{U}$) give rise to a cq-wiretap channel $\mathcal{M}:\mathcal{U}\rightarrow(BB^{\prime})^{n}(EE^{\prime})^{n}$, mapping $u\in\mathcal{U}$ to $\mathcal{M}(u)=(\mathcal{N}^{\otimes n}\!\circ\mathcal{E}_{u})\zeta^{\otimes n}$. We can thus apply the coding theorem and weak converse from [10] to get the following expression for the assisted private capacity:
@@
V Discussion
We have presented a coding strategy for the quantum wiretap channel assisted by a general quantum correlation, which shows a systematic improvement of the private communication rate over the unassisted private capacity of the channel due to the exploitation of the correlation. Our Theorem 5 also directly generalises the main result of [12], since classical channels are a special case of cptp maps and a classical correlation between random variables $X$, $Y$ and $Z$ is naturally represented by the diagonal state $\zeta^{A^{\prime}B^{\prime}E^{\prime}}=\sum_{xyz}P_{XYZ}(xyz)\left|x\right\rangle\!\left\langle x\right|^{A^{\prime}}\otimes\left|y\right\rangle\!\left\langle y\right|^{B^{\prime}}\otimes\left|z\right\rangle\!\left\langle z\right|^{E^{\prime}}$. As the proposed code uses a method derived from the (quantum) Gel’fand-Pinsker wiretap channel with side information, it is also always at least as good as, or superior to, the “trivial” incorporation of the correlation resource into an augmented wiretap channel. However, due to the general need for regularisation, both our main result (Theorem 5) and the “trivial” achievable rate [Eq. (5)] lead to multi-letter formulas for the same quantity $P(\mathcal{N},\zeta)$.
@@ V Discussion
Now this achievability bound and expression for the assisted capacity correspond exactly to Eq. (3), and interestingly it provides an instance where Theorem 5 is better, at least on the single-letter level (naturally, in the regularisation both expression result in the same number). Namely, in [22] a certain duality relation was observed between the dense coding advantage and the so-called entanglement of purification $E_{P}$ [28], which extends to the regularisations of the two quantities:
@@
It remains unknown whether $\Delta=\Delta^{(\infty)}$ and $E_{P}=E_{P}^{(\infty)}$, but in [11] numerical evidence to the contrary was provided for the entanglement of purification. Showing $E_{P}(\rho)>E_{P}^{(\infty)}(\rho)$ for a certain state in [11] required calculating the l.h.s. numerically, and finding a new upper bound for the r.h.s. via an asymptotic protocol using the covering lemma. Concretely, it was shown for an ensemble $\{q(u),\rho_{u}^{CD}\}$ with $\rho=\sum_{u}q(u)\rho_{u}$ that
@@
where $\gamma^{UCD}=\sum_{u}q(u)\left|u\right\rangle\!\left\langle u\right|^{U}\otimes\rho_{u}^{CD}$ is the cq-state of the ensemble. Using the duality relation from [22] to translate this protocol to $\Delta^{(\infty)}$, it corresponds to the achievable rate from Theorem 5 in the case $\mathcal{N}=\operatorname{id}_{A}$ and $\zeta^{A^{\prime}B^{\prime}}$.
@@ Acknowledgments
MLC was supported by the German Research Foundation (DFG) under the Walter Benjamin Fellowship CA-2779/1-1, and in part by the ESA SATNEX V programme (project 4000130962/20/NL/NL/FE). Both authors acknowledge support by the Spanish MICIN (project PID2022-141283NB-I00) with the support of FEDER funds. AW is furthermore supported by the European Commission QuantERA grant ExTRaQT (Spanish MICIN project PCI2022-132965), the Spanish MICIN with funding from European Union NextGenerationEU (PRTR-C17.I1) and the Generalitat de Catalunya, by the Alexander von Humboldt Foundation, and the Institute for Advanced Study of the Technical University Munich.